Sodium hydroxide reacts with carbon dioxide according to the equation: NaOH(s) + CO2(g) → Na2CO3(s) + H2O(l). a) Which is the "limiting reagent" when 7.4 g of NaOH and 0.44 g of CO2 are allowed to react? b) How many grams of Na2CO3 are expected to be produced ? c) What would be the percent yield if 0.90 g of Na2CO3 are obtained?
1
Expert's answer
2020-11-17T10:37:46-0500

2NaOH(s)+CO2(g)=Na2CO3(s)+H2O(l)

Moles of NaOH in 7.4g

Moles=Mass/RMM

=7.4g/40gmol-1

=0.185moles

Moles of CO2 in 0.44g =0.44g/44gmol-1

=0.01moles

If 0.01 moles of CO2 were to react then, 2(0.01) moles NaOH was to be consumed

=0.02 moles of NaOH

If 0.185 moles of NaOH reacted then, 0.185moles/2 CO2 reacted

=0.0925moles of CO2

Since the number of moles of CO2 were less than 0.0925 moles, it means CO2 was the limiting reagent as all of it, CO2, reacted while (0.185-0.02)g = 0.165g of NaOH remained in excess


Theoretical mass of Na2CO3 that was to be formed

Theoretical Moles of Na2CO3= 0.01moles

mass=0.01g*106gmol-1

=1.06g


%yield of Na2CO3= (0.90g/1.06g)*100%

=84.91%